# PhoenixOS: Concurrent OS-level GPU Checkpoint and Restore with Validated Speculation

### v3

@@ Introduction
What is and why concurrent OS-level checkpoint and restore (C/R).    OS-level checkpoint snapshots the execution of a running process as an image, which the OS can then use to restore the execution of the process. It is a foundational OS primitive: cluster management systems rely on it to migrate tenant jobs (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/osdi/XiaoBRSKHPPZZYZ18, ; cuda-migration, )—by first checkpointing the image to the target machine, then restoring from it. Serverless systems leverage OS-level restore to quickly start new processes (serverless-gpu-ali, ; aws-snapstart, ; DBLP:conf/asplos/DuYXZYQWC20, ). Cloud providers periodically checkpoint computing jobs to provide fault tolerance in case of failures (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/sosp/Wu0M023, ).
A key feature of OS-level checkpoint and restore is that it is transparency to the processes, making it irreplaceable for functionality and efficiency. For functionality, consider cloud providers that need to migrate tenant jobs to improve cluster utilization and for other purposes (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/osdi/XiaoBRSKHPPZZYZ18, ; cuda-migration, ). Since these jobs operate as black boxes to the providers (cuda-migration, ), migration can only be performed at the OS level. For efficiency, consider serverless platforms that need to quickly start GPU processes (ali-serverless-gpu, ). These processes have PyTorch runtime states (e.g., compiler cache) that are tightly intertwined with OS information. Recreating these states from scratch takes seconds; thankfully, OS-level C/R can significantly reduce this overhead to milliseconds (DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/Ustiugov0KBG21, ).
Additionally, OS-level C/R can also provide superior usability (DBLP:journals/corr/abs-2202-07848, ; microsoft-forge, ), which is why major cloud providers (e.g., Microsoft Forge (microsoft-forge, ; DBLP:conf/eurosys/GuptaKKVGKRS24, )) employ it for fault tolerance. The key rationale is that performing efficient and correct fault tolerance through user-level C/R is challenging due to the broad optimization space and quickly evolving workloads (DBLP:conf/sosp/WangJZZFNW23, ; DBLP:conf/fast/MohanPC21, ; DBLP:journals/corr/abs-2407-20143, ). For instance, the leading LLM training framework Megatron (megatron-checkpoint, ) has recently integrated concurrent checkpoint optimizations proposed three years ago (DBLP:conf/fast/MohanPC21, ), with the checkpoint code now comprising a quarter of its codebase. Meanwhile, many emerging training frameworks, such as those for reinforcement learning (gym, ), still lack these features. OS-level C/R avoids the developers’ burden of implementing and optimizing their own fault tolerance mechanisms.
@@
Table 1. An overview of the information (Info.) required for correct concurrent checkpoint and restore, and how PhOS traces them.
@@
Retrofitting CPU-based concurrent C/R techniques on GPUs with our validated speculation—although ensuring correctness—still suffers from performance issues if not considering GPU features carefully. First, during concurrent checkpoint, the interference between CPU and GPU checkpoint, as well as between application data transfers and checkpoint transfers, can significantly impact the efficiency of the concurrent checkpoint. To mitigate these interferences, we design a coordinated and prioritized checkpoint mechanism (§5). Second, before allowing a concurrent restore, we must create the proper GPU execution environment, which incurs overhead comparable to that of data restoration itself. We propose using a GPU context pool to bypass the context creation to fully unleash the power of concurrent restore (§6).
@@
Demonstrations.  PhOS can C/R unmodified GPU applications (including multi-GPU processes) on NVIDIA GPUs. We chose NVIDIA GPUs due to their prevalent adoption, despite higher implementation complexity. Nevertheless, our design also generalizes to other GPUs and accelerators, because they follow the same execution model of NVIDIA GPUs (§2.1). On A800 GPU servers, when compared to state-of-the-art systems—Singularity (DBLP:journals/corr/abs-2202-07848, ) and NVIDIA’s official OS-level C/R (cuda-checkpoint) (nvidia-new-ckpt-driver-repo, ), PhOS reduces the application stall caused by checkpoint and restore by up to 160%. More importantly, PhOS brings orders of magnitude end-to-end application performance improvements. Notably, compared to Singularity, it reduces the wasted GPU time caused by fault tolerance by 76% when training Llama3-13B on multiple GPUs, it reduces the downtime caused by migrating a Llama2-13B inference job from 9.8 to 2.3 seconds, and it can cold-start a Llama2-13B inference job within only 622 milliseconds, which is 124–450% and 114–342% faster than cuda-checkpoint and Singularity, respectively.
@@
  • The first set of OS-level concurrent checkpoint and restore protocols for GPUs: §4.2, §4.3, and §6.
  • PhOS is the first OS-level concurrent checkpoint and restore system for GPUs. It addresses several technical challenges to realize the concurrent protocols (§5 and §6), and it significantly enhances various end-to-end application performance (§8).
@@
When a process starts, the GPU driver creates execution context, e.g., CUcontext (cuda-context, ), involving compiling and loading the kernel binaries, configuring the GPU virtual memory and others (Mai2023HONEYCOMB, ). Afterward, processes use GPU driver or toolkit APIs (we term GPU API in this paper) like CUDA (cuda-driver, ), illustrated in Figure 1 (a), to trigger computations on GPUs. For example, the process can launch kernels via cudaLaunchKernel and copy CPU data to GPU (and vice versa) using cudaMemcpy. PhOS intercepts GPU APIs for realizing checkpointing and restoration, thus supporting all GPU applications at the OS-level.
@@
Basic checkpoint and restore.    A checkpoint captures a process’s execution state at a specific time in an image (termed as checkpoint). The OS can then use this checkpoint to recreate (restore) the process. A checkpoint includes two data categories: data state (data in the CPU virtual memory and GPU buffers) and control state (e.g., CPU registers that store the program counter, PC). The checkpointed CPU states also include kernel objects like network connections (criu-tcp, ; Aurora, ).
A common approach for implementing GPU checkpointing is by first quiescing process execution (both CPU and GPU), i.e., stops the CPU execution and waiting for all in-flight GPU kernels and communications to finish, then copying all aforementioned data to the checkpoint (nvidia-new-ckpt-driver-repo, ; DBLP:journals/corr/abs-2202-07848, ). Figure 1 (b) illustrates this. For restoration, the OS first loads data state into CPU memory and GPU buffers, then restores control state to resume execution.
@@
Our command-line tool (❶) enables users to checkpoint running GPU processes, migrate them across machines, or restore processes from checkpoints with the process ID (\$PID). Though PhOS does not require application cooperation, we found that it can benefit from choosing a proper checkpoint time (see §8.4), so we also provide an SDK for applications to control the timing with only six lines of additional code (Details are left in §A.2 due to space limitation).
@@
Our backend (❸) uses CRIU (criu, ) to checkpoint and restore CPU process states, and our efficient GPU engine (see §5) for GPU states. Choosing CRIU allows PhOS to correctly checkpoint and restore complex states like network connections. However, CRIU may not be the most efficient solution for CPU checkpointing. Our current implementation overlaps the CRIU checkpointing with our GPU checkpointing but a faster CPU checkpointing would make PhOS easier to implement.
@@
PhOS command-line tool uses inter-process calls (IPC) to communicate with the frontend. On the other hand, the frontend and backend can communicate either via IPC or local function calls (LFC). We prefer LFC whenever feasible by linking the frontend and backend in the same library. However, IPC is required for cases when the applications require accessing the context pool, e.g., for fast restore (see §7). This is because the context must be pre-created in PhOS daemon to bypass the costly context creation during restore (see Figure 2). This comes at the cost of extra GPU API calling overhead during application running (excluding the restore), e.g., we observed a maximum of 9% IPC overhead for typical AI applications after applying recent works in accelerating calling GPU APIs with IPC (DBLP:journals/corr/abs-2401-13354, ; DBLP:journals/corr/abs-2306-03622, ). We believe it is a proper trade-off as the restore time dominates the overall application execution time in the required scenarios (see §7).
@@
Figure 4. An illustration of correctness guarantee of concurrent checkpointing enabled by different protocols of PhOS.
@@
Both protocols require tracing the process’s write set, i.e., which bytes are written during concurrent checkpoint. For CPU writes, we follow the traditional approach to use permission bits in the page table, i.e., write protection for copy-on-write and dirty bits to trace at the page granularity. We omit the detailed CPU-side description since it is well studied. Unfortunately, modern GPUs lack these support, e.g., no hardware dirty bits (DBLP:conf/uss/0002Z024, ). Therefore, before delving into our protocols, we first propose a software-based approach to trace GPU writes.
@@
Tracing with speculation on GPU API arguments.    All operations capable of performing GPU writes are initiated by calling the GPU API, so we intercept these calls to trace the write set. These APIs fall into four categories:
@@
    Issue memory move operations, e.g., cudaMemcpy (dst, src, count, ...) . They write the CPU data (src) to the GPU memory (dst).
@@
    Issue communication kernels, e.g., ncclBroadcast (sendbuffer, recvbuffer, count,...). They write the GPU buffers with the content sent through the network.
@@
    Issue computation kernels with well-defined semantics, e.g., cublasSgemm(...,A,B,...,C,...). They perform a computation and update the result buffers (C), whose read and write semantics are know in the specifications (cuBLAS, ).
@@
    Issue opaque computation kernels, e.g., cudaLaunch -Kernel (func, ... , args, ...) . They launch kernels written by the developers or just-in-time compiled during runtime, so the reads and writes are unknown to the OS.
For types 1–3 APIs, PhOS directly uses their specifications to trace writes. For instance, the specification of cudaMemcpy tells us that executing this API will modify the GPU memory from dst to dst + count. Empirically, we found over 50% invocations are such APIs (§8.2).
@@
Opaque kernels are trickier because the OS only knows the kernel binary and the arguments invoking it. Our observation is that each GPU kernel represents a clear and straightforward computational purpose whose accessed data addresses are directly encoded in the OS-known arguments calling the kernel. As shown in Figure 5, the data (output) written by the kernel is the second argument. Since the OS knows all the buffers allocated by the process by intercepting the GPU memory allocation calls, line 1 (cudaMalloc), and it knows each kernel’s function signature (cuda-signature, ), we can systematically compare the arguments with the allocated buffers to speculate the buffer written by each kernel. The intercepting overhead is negligible: it only executes a few data structures manipulations atop of existing GPU memory allocation APIs.
@@
When PhOS intercepts a kernel invocation during checkpoint (❸), it invokes its corresponding instrumented twin kernel. If the kernel reports validation failures, we will execute fallback protocols (described in §4.2 and §4.3) to ensure correctness. The overhead of running the instrumented kernels is also manageable (§8.2) and they are not invoked without checkpoint (❷).
Discussion: the granularity of tracing.    Ideally, the trace should be as fine-grained as possible to avoid protocol overhead for ensuring correctness, e.g., avoid excessive copy-on-write. For non-opaque kernels, we can precisely track the bytes written. However, our speculation-based tracing can only detect writes at the buffer-level, which may result in over-tracing. For example, when a kernel only writes a small part of the buffer, we may treat the entire buffer as written. Luckily, such over-tracing is rare especially for recent GPU applications like AI jobs, because (1) AI frameworks (e.g., PyTorch) use a fine-grained buffer allocation, e.g., one for each tensor (pytorch-mem, ), and (2) the opaque kernels typically write the entire buffer (tensor). See §8.5 for the analysis.
@@
Suppose a checkpoint starts at time t1. Our CoW protocol guarantees the final checkpoint matches a stop-the-world checkpoint at t1, while allowing concurrent application execution. At a high level, the inconsistency can only happen during execution if the GPU writes to an uncheckpointed buffer. To fix the issue, before we allow such writes to execute, we first copy the targeted data to another place, e.g., a free on-device buffer; then redirect application writes to a new buffer. Thus, the concurrent checkpoint will not see the new writes.
@@
Second, we only reserve a small GPU memory (up to 2 GB) for the copy-on-write, and the concurrent checkpoint can still work even if no extra memory is used. If we lack sufficient GPU memory for the copy-on-write, we block the kernel until free memory is available (K2 in Figure 7) or the written buffer has been checkpointed. This causes a little checkpoint stall but we found it acceptable: PhOS does not require a large on-device buffer for high performance because once a copy-on-write is done, its original buffer can be immediately released. §8.3 analyzes the impact of GPU memory reserved for CoW on the checkpointing performance.
Finally, if the validator detects a mis-speculation by reporting a write to a buffer not traced, we will check whether it affects correctness and react if necessary. If PhOS has checkpointed the buffer before kernel execution, no further action is required. Otherwise, we discard the current checkpoint and retry with a stop-of-the-world approach for liveness. Note that because we don’t encounter speculation failures for all major GPU applications, we adopted a simple retry strategy and leave a more advanced one as our future work.
Correctness.    Our checkpoint is the same as the resulting checkpoint from a stop-of-the-world checkpoint at the start checkpoint time. First, our quiesce phase is the same as the stop checkpoint. So before the copy phase, our initial process states are the same as a stop checkpoint. Second, during our concurrent copy, we only copy the states not modified by concurrent execution, because all writes are isolated by copy-on-write. As the stop checkpoint also faithfully checkpoints the same state, our checkpoint is the same as its checkpoint.
Discussion: speculation vs. validation.    Although we can, in principle, avoid speculation by instrumenting the kernels to obtain their accessed buffers and do the copy-on-write, as we did in validation, we still choose speculation for two reasons. First, speculation allows us to acquire buffer information in advance, which is necessary for CoW especially in case of insufficient on-device buffer. Besides, our concurrent restore protocol (see §6) also relies on knowing the accessed buffers in advance. Second, speculation then validation has fewer overheads than obtaining the buffers without speculation, because in the common case, the validation will only pass a flag indicating whether the validation passes.
@@
Reducing dirty buffers via coordinated GPU and CPU checkpoint.    In the soft recopy protocol, the time window of the copy is critical to the number of dirty buffers generated: the longer the copy, the more likely the buffers are becoming dirty. If we copy the CPU and GPU data without coordination, the overall time window is the size of checkpoint image divided by the copy bandwidth.
In GPU processes, we found GPU writes data more frequently than CPU writes, so it is beneficial to first copy the CPU data instead of copying the GPU data, reducing the GPU time window to the size of the GPU data divided by the copy bandwidth, as shown in Figure 9 (b). Thus, we implement a coordinated checkpoint in our checkpoint engine: we first copy the CPU data, then the GPU data.
Prioritized application PCIe transfer.    When we completed our first concurrent checkpoint implementation (for both protocols), we found the process still meets significant stalls during the checkpoint. By profiling the GPU execution, we found the stall is not caused by our protocols, but by the saturated DMA engine caused by GPU checkpoint. Specifically, GPUs have a limited number of DMA engines (bakitademystifying, ) shared between PhOS and applications. With concurrent checkpoint, the bulk copy nature of the checkpoint can easily saturate the DMA engines, causing application kernels to wait.
To this end, we adopt two techniques to avoid application starvation—the goal is to prioritize application DMA transfer over the checkpoint, observing that applications infrequently use the DMA but such usages are on their critical path. First, we intercept all process’s DMA-related API calls (e.g., cudaMemcpy) to detect application PCIe transfer. The interception overhead is negligible (<1%), because the APIs are typically called asynchronously so our interception is not on the application’s critical path (DBLP:journals/corr/abs-2401-13354, ; DBLP:journals/corr/abs-2306-03622, ). Based on the detection, we implemented a preemptible checkpoint copy mechanism for the application data copy: for each GPU buffer, we copy it to the checkpoint in small 4 MB chunks. After copying one chunk, we check whether there is ongoing or pending application transfer. If so, we pause the checkpoint copy and let the pending application transfers execute first.
@@
Handling mis-speculation.    Like writes, we also instrument a validator to ensure correctness under speculation failures. If the speculation is wrong, the validator will notify PhOS to handle. However, the mis-speculation handling of reads is more complex than writes, because the kernel may have accessed inconsistent GPU states caused by a partially restored buffer. To ensure correctness, we must roll back the GPU states to a correct one. Our current solution rolls back to the initial state from the checkpoint and then performs a stop-of-the-world restore for liveness. We choose this simple solution because we have not met speculation failures in common GPU applications. While a more efficient solution could involve isolating partial updates using our soft copy-on-write protocol, we defer the detailed design to future work.

### v4

@@ Abstract.
GPU checkpoint and restore, concurrent checkpoint and restore, validated speculation
@@ Introduction
What is concurrent OS-level checkpoint and restore (C/R) and why it matters.    OS-level checkpoint snapshots the execution of a running process as an image, which the OS can later use to restore the process. This is a foundational OS primitive with many key applications: cluster management systems rely on it to migrate tenant jobs (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/osdi/XiaoBRSKHPPZZYZ18, ; cuda-migration, ) by first checkpointing the image to the target machine and then restoring from it. Serverless systems leverage OS-level restore to launch new processes quickly (serverless-gpu-ali, ; aws-snapstart, ; DBLP:conf/asplos/DuYXZYQWC20, ). Cloud providers implement periodic checkpointing to provide fault tolerance against failures (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/sosp/Wu0M023, ).
A key feature of OS-level checkpoint and restore is that it is transparent to processes, making it irreplaceable for both functionality and efficiency. For functionality, cloud providers need migrate tenant jobs to improve cluster utilization and other purposes (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/osdi/XiaoBRSKHPPZZYZ18, ; cuda-migration, ). Since these jobs operate as black boxes to the providers (cuda-migration, ), migration can only be performed at the OS level. For efficiency, serverless platforms need to launch GPU processes quickly (ali-serverless-gpu, ). These processes contain PyTorch runtime states (e.g., compiler cache) that are tightly intertwined with OS information. Rebuilding these states from scratch takes seconds; thankfully, OS-level C/R can significantly reduce this overhead to milliseconds (DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/Ustiugov0KBG21, ).
Additionally, OS-level C/R can also provide superior usability (DBLP:journals/corr/abs-2202-07848, ; microsoft-forge, ), which is why major cloud providers (e.g., Microsoft Forge (microsoft-forge, ; DBLP:conf/eurosys/GuptaKKVGKRS24, )) employ it for fault tolerance. The key rationale is that performing efficient and correct fault tolerance through user-level C/R is challenging due to the broad optimization space and quickly evolving workloads (DBLP:conf/sosp/WangJZZFNW23, ; DBLP:conf/fast/MohanPC21, ; DBLP:journals/corr/abs-2407-20143, ). For instance, the leading LLM training framework Megatron (megatron-checkpoint, ) has recently integrated concurrent checkpoint optimizations proposed three years ago (DBLP:conf/fast/MohanPC21, ), with the checkpoint code now comprising a quarter of its codebase. Meanwhile, many emerging training frameworks, such as those for reinforcement learning (gym, ), still lack these features. OS-level C/R does not impose any burden on developers for implementing and optimizing their own fault tolerance mechanisms.
@@
Table 1. An overview of information required for correct concurrent checkpoint and restore, and how PhOS traces them.
@@
Retrofitting CPU-based concurrent C/R techniques on GPUs with our validated speculation—although ensuring correctness—still suffers from performance issues if GPU features are not considered. First, during concurrent checkpointing, the interference between CPU and GPU checkpointing, as well as between application data transfers and checkpoint transfers, can significantly impact the efficiency of the concurrent checkpointing. To mitigate this interference, we design a coordinated and prioritized checkpoint mechanism (§5). Second, before allowing a concurrent restore, we must create the proper GPU execution environment, which incurs overhead comparable to that of data restoration itself. We propose using a GPU context pool to bypass the context creation to fully unleash the power of concurrent restore (§6).
@@
Demonstrations.  PhOS can checkpoint and restore unmodified GPU applications (including multi-GPU processes) on NVIDIA GPUs. We selected NVIDIA GPUs for their widespread adoption, despite the greater implementation challenges they present. Our design also generalizes to other GPUs and accelerators that follow the same execution model of NVIDIA GPUs (§2.1). When evaluated on A800 GPU servers and compared to state-of-the-art systems—Singularity (DBLP:journals/corr/abs-2202-07848, ) and NVIDIA’s official checkpoint and restore utility (cuda-checkpoint) (nvidia-new-ckpt-driver-repo, ), PhOS reduces application stall during checkpoint and restore by up to 160%. More importantly, PhOS delivers orders-of-magnitude improvements in end-to-end application performance. Compared to Singularity, PhOS reduces wasted GPU time in fault tolerance by 76% when training Llama2-13B on multiple GPUs. It also decreases migration downtime for a Llama2-13B inference job from 9.8 to 2.3 seconds, Additionally, PhOS can launch a new Llama2-13B inference job in just 622 milliseconds by avoiding cold start costs, which is 114–342% and 124–450% faster than Singularity and cuda-checkpoint, respectively.
@@
  • The first set of OS-level concurrent checkpoint and restore protocols for GPUs (§4.2, §4.3, and §6).
  • PhOS, the first OS-level concurrent checkpoint and restore system for GPUs that addresses critical technical challenges to realize these concurrent protocols (§5 and §6), significantly improving end-to-end application performance (§8).
@@
When a process starts, the GPU driver creates an execution context, e.g., CUcontext (cuda-context, ), involving compiling and loading the kernel binaries, configuring the GPU virtual memory and others (Mai2023HONEYCOMB, ). Afterward, processes use a GPU driver or toolkit APIs (we term GPU API in this paper) like CUDA (cuda-driver, ), illustrated in Figure 1(a), to trigger computations on GPUs. For example, the process can launch kernels via cudaLaunchKernel and copy CPU data to GPU (and vice versa) using cudaMemcpy. PhOS intercepts GPU APIs for realizing checkpointing and restoration, thus supporting all GPU applications at the OS-level.
@@
Basic checkpoint and restore.    A checkpoint captures a process’s execution state at a specific time in an image (termed as checkpoint). The OS can then use this checkpoint to recreate the process (termed as restore). A checkpoint contains two types of data: data state (data in the CPU memory and GPU buffers) and control state (e.g., CPU registers that store the program counter, PC). Checkpointed CPU states also include kernel objects like network connections (criu-tcp, ; Aurora, ).
A common approach for implementing GPU checkpointing first quiesces (both CPU and GPU) process execution—stopping CPU execution and waiting for all in-flight GPU kernels and communications to complete—and then copies all execution state to the checkpoint (nvidia-new-ckpt-driver-repo, ; DBLP:journals/corr/abs-2202-07848, ), as illustrated in Figure 1(b). For restoration, the OS first loads data state into CPU memory and GPU buffers, then restores control state to resume execution.
@@
Our command-line tool (❶) enables users to checkpoint running GPU processes, migrate them across machines, or restore processes from checkpoints with the process ID (\$PID). Although PhOS does not require application cooperation, we found that it can benefit from choosing a proper checkpoint time (see §8.4). We therefore provide an SDK for applications to control checkpoint timing with just five lines of additional code (more details in §A.2).
@@
Our backend (❸) uses CRIU (criu, ) to checkpoint and restore CPU process states, and our efficient GPU engine (see §5) for GPU states. Choosing CRIU allows PhOS to correctly checkpoint and restore complex states like network connections. However, CRIU may not be the most efficient solution for CPU checkpointing. Our current implementation overlaps CRIU checkpointing with our GPU checkpointing but a faster CPU checkpointing mechanism would make PhOS easier to implement.
@@
PhOS command-line tool uses inter-process calls (IPC) to communicate with the frontend. On the other hand, the frontend and backend can communicate either via IPC or local function calls (LFC). We prefer LFC whenever feasible by linking the frontend and backend in the same library. However, IPC is required for cases when the applications require accessing the context pool, e.g., for fast restore (see §7). This is because the context must be pre-created in PhOS daemon to bypass the costly context creation during restore (see Figure 2). This comes at the cost of extra GPU API calling overhead when the application executes (after the restore), e.g., we observed a maximum of 9% IPC overhead for typical AI applications after applying recent works in accelerating calling GPU APIs with IPC (DBLP:journals/corr/abs-2401-13354, ; DBLP:journals/corr/abs-2306-03622, ). We believe this is a reasonable trade-off as the restore time dominates the overall application execution time in the required scenarios (see §7).
@@
Figure 4. An illustration of the correctness guarantee of concurrent checkpointing enabled by different protocols of PhOS.
@@
Both protocols require tracing the process’s write set, i.e., which bytes are written during the concurrent checkpoint. For CPU writes, we follow the traditional approach to use permission bits in the page table, i.e., write protection for copy-on-write and dirty bits to trace at the page granularity. We omit the detailed CPU-side description since it is well studied. Unfortunately, modern GPUs lack these supports, e.g., no hardware dirty bits (DBLP:conf/uss/0002Z024, ). Therefore, before delving into our protocols, we first propose a software-based approach to trace GPU writes.
@@
Tracing with speculation on GPU API arguments.    All operations that can perform GPU writes are initiated through GPU API calls, so we intercept these calls to trace the write set. These APIs fall into four categories:
@@
    Issue memory move operations, e.g., cudaMemcpy(dst,
    src,count,..)
    . They write the CPU data (src) to the GPU memory (dst).
@@
    Issue communication kernels, e.g., ncclBroadcast
    (sendbuffer,recvbuffer,count,..)
    . They write the GPU buffers with the content sent through the network.
@@
    Issue computation kernels with well-defined semantics, e.g., cublasSgemm(..,A,B,..,C,..). They perform a computation and update the result buffers (C), whose read and write semantics are known in the specifications (cuBLAS, ).
@@
    Issue opaque computation kernels, e.g., cudaLaunch-
    Kernel(func,..,args,..)
    . They launch kernels written by the developers or just-in-time compiled during runtime, so the reads and writes are unknown to the OS.
For types 1–3 APIs, PhOS directly uses their specifications to trace writes. For instance, the specification of cudaMemcpy indicates that executing this API will modify GPU memory from dst to dst+count. Our empirical analysis shows that over 50% of invocations are these types of APIs (§8.2).
@@
Opaque kernels are trickier because the OS only knows the kernel binary and the arguments invoking it. Our observation is that each GPU kernel represents a clear and straightforward computational purpose whose accessed data addresses are directly encoded in the OS-known arguments calling the kernel. As shown in Figure 5, the data (output) written by the kernel is the second argument. Since the OS knows all the buffers allocated by the process by intercepting the GPU memory allocation calls, line 1 (cudaMalloc), and it knows each kernel’s function signature (cuda-signature, ), we can systematically compare the arguments with the allocated buffers to speculate the buffer written by each kernel. The interception overhead is negligible: it only manipulates a few data structures before calling GPU memory allocation APIs.
@@
When PhOS intercepts a kernel invocation during checkpoint (❸), it invokes its corresponding instrumented twin kernel. If the kernel reports validation failures, we will execute fallback protocols (described in §4.2 and §4.3) to ensure correctness. The overhead of running the instrumented kernels is also manageable (§8.2) and they are not invoked without a checkpoint (❷).
Figure 6. PhOS validation workflow.
Discussion: the granularity of tracing.    Ideally, the trace should be as fine-grained as possible to avoid protocol overhead for ensuring correctness, e.g., avoid excessive copy-on-write. For non-opaque kernels, we can precisely track the bytes written. However, our speculation-based tracing can only detect writes at the buffer-level, which may result in over-tracing. For example, when a kernel only writes a small part of the buffer, we may treat the entire buffer as written. Luckily, such over-tracing is rare, especially for recent GPU applications like AI jobs, because (1) AI frameworks (e.g., PyTorch) use a fine-grained buffer allocation, e.g., one for each tensor (pytorch-mem, ), and (2) the opaque kernels typically write the entire buffer (tensor). See §8.5 for the analysis.
@@
Suppose a checkpoint starts at time t1. Our CoW protocol guarantees the final checkpoint matches a stop-the-world checkpoint at t1, while allowing concurrent application execution. The inconsistency can only happen when the GPU writes to an uncheckpointed buffer during execution. To prevent this, before allowing such writes to execute, we first copy the targeted data to another place (e.g., a free on-device buffer) and then redirect application writes to a new buffer. Thus, the concurrent checkpoint will not see the new writes.
@@
Second, we only reserve a small amount of GPU memory (up to 2 GB) for the copy-on-write, and the concurrent checkpoint can still work even if no extra memory is used. If we lack sufficient GPU memory for the copy-on-write, we block the kernel until free memory is available (K2 in Figure 7) or the written buffer has been checkpointed. This causes a short checkpoint stall but we have found it acceptable: PhOS does not require a large on-device buffer for high performance because once a copy-on-write is done, its original buffer can be immediately released. §8.3 analyzes the impact of GPU memory reserved for CoW on the checkpointing performance.
Finally, if the validator detects a mis-speculation by reporting a write to a buffer not traced, we will check whether it affects correctness and react if necessary. If PhOS has checkpointed the buffer before kernel execution, no further action is required. Otherwise, we discard the current checkpoint and retry with a stop-the-world approach for liveness. Note that because we don’t encounter speculation failures for all major GPU applications, we adopted a simple retry strategy and leave a more advanced one as our future work.
Correctness.    Our checkpoint is the same as the resulting checkpoint from a stop-the-world checkpoint at the start checkpoint time. First, our quiesce phase is the same as the stop checkpoint. So before the copy phase, our initial process states are the same as a stop checkpoint. Second, during our concurrent copy, we only copy the states not modified by concurrent execution, because all writes are isolated by copy-on-write. As the stop checkpoint also faithfully checkpoints the same state, our checkpoint is the same as its checkpoint.
Discussion: speculation vs. validation.    Although we can, in principle, avoid speculation by instrumenting the kernels to obtain their accessed buffers and do the copy-on-write, as we did in validation, we still choose speculation for two reasons. First, speculation allows us to acquire buffer information in advance, which is necessary for CoW especially in case of insufficient on-device buffer. Also, our concurrent restore protocol (see §6) relies on knowing the accessed buffers in advance. Second, speculation then validation has fewer overheads than obtaining the buffers without speculation, because in the common case, the validation will only pass a flag indicating whether the validation passes.
@@
Reducing dirty buffers via coordinated GPU and CPU checkpoint.    In the soft recopy protocol, the time window of the copy is critical to the number of dirty buffers generated: the longer the copy, the more likely the buffers will become dirty. If we copy the CPU and GPU data without coordination, the overall time window is the size of checkpoint image divided by the copy bandwidth.
In GPU processes, we found GPU writes are more frequent than CPU writes, so it is beneficial to first copy the CPU data instead of copying the GPU data, reducing the GPU time window to the size of the GPU data divided by the copy bandwidth, as shown in Figure 9(b). Thus, we implement a coordinated checkpoint in our checkpoint engine: we first copy the CPU data, then the GPU data.
Prioritized application PCIe transfer.    When we completed our first concurrent checkpoint implementation (for both protocols), we found the process still has significant stalls during the checkpoint. By profiling the GPU execution, we found the stall is not caused by our protocols, but by the saturated DMA engine caused by GPU checkpointing. Specifically, GPUs have a limited number of DMA engines (bakitademystifying, ) shared between PhOS and applications. With concurrent checkpointing, the bulk copy nature of the checkpointing can easily saturate the DMA engines, causing application kernels to wait.
To this end, we adopt two techniques to avoid application starvation—the goal is to prioritize application DMA transfer over the checkpoint, observing that applications infrequently use DMA but DMA operations are on the critical path. First, we intercept all process’s DMA-related API calls (e.g., cudaMemcpy) to detect application PCIe transfer. The interception overhead is negligible (<1%), because the APIs are typically called asynchronously so our interception is not on the application’s critical path (DBLP:journals/corr/abs-2401-13354, ; DBLP:journals/corr/abs-2306-03622, ). When the DMA operation is detected, we implemented a preemptible checkpoint copy mechanism: for each GPU buffer, we copy it to the checkpoint in small 4 MB chunks. After copying one chunk, we check whether there is ongoing or pending application transfer. If so, we pause the checkpoint copy and let the pending application transfers execute first.
@@
Handling mis-speculation.    Like writes, we also instrument a validator to ensure correctness under speculation failures. If the speculation is wrong, the validator will notify PhOS to handle. However, the mis-speculation handling for reads is more complex than for writes, because the kernel may have accessed inconsistent GPU states caused by a partially restored buffer. To ensure correctness, we must roll back the GPU states to a correct one. Our current solution rolls back to the initial state from the checkpoint and then performs a stop-the-world restore for liveness. We choose this simple solution because we have not met speculation failures in common GPU applications. While a more efficient solution could involve isolating partial updates using our soft copy-on-write protocol, we defer the detailed design to future work.